# Identities are not Interchangeable: The Problem of Overgeneralization in Fair Machine Learning

### v1

@@ Abstract.
journalyear: 2025copyright: noneconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732033isbn: 979-8-4007-1482-5/2025/06ccs: Social and professional topics User characteristicsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Central to most fair machine learning algorithms and measurements are demographic axes, and the groups within the axis. For example, fairness evaluations will measure the difference in outcomes across the demographic axis of gender for the groups of men and women. However, which identity axes or groups usually do not matter to the algorithm or measurement, as the axis is generally left an open variable. In this work, we argue for the importance of specificity in demographic axis. In other words, you cannot necessarily build or measure the fairness of a machine learning system which is fair with respect to race the same way you can for one with respect to gender.
Fair machine learning has inherited the propensity from machine learning to seek abstractions and generalization, prioritizing methods which are domain-agnostic (Birhane et al., 2022). The ideal model architecture is one which works well across many data distributions. This mentality has led to methods that, for example, treat discrimination as a disparity that occurs between any two social groups. This is not always a bad thing. The ease of implementation often determines whether a quantity is measured at all (Wang et al., 2024), so shoehorning a measure for an overlooked axis (e.g., disability discrimination) into an existing pipeline (e.g., for measuring racial discrimination), can bring more attention than if axis-specific pipelines were to be established in each setting. Similarly, once pipelines for multi-group fairness exist, it seems that so too do those for intersectionality if it can be massaged into the same technical format (Wang et al., 2022; Kong, 2022). Because of the technical convenience of one-size-fits-all fairness, we have generally desired one “fair” algorithm and one “fair” definition across all domains. There is increasingly recognition for domain-specific considerations that will affect the relevant definitions of fairness. However, an often-overlooked aspect is the variability in the axis of identity being targeted. For instance, whether racism in fair machine learning be studied and addressed differently from sexism in fair machine learning. A common refrain we see in the problem formulation of technical machine learning papers on group fairness is a phrase like “$A$ represents the protected attribute (e.g., race or gender). $A=1$ is the privileged group, and $A\neq 1$ is the unprivileged group” (Mehrabi et al., 2021; Pessach and Shmueli, 2022). But does it matter whether $A$ is race or gender? What about age, or disability?
@@ 2.2. Differences
In considering various forms of oppression, there are debates about the primacy and causal relationships amongst them. Karl Marx believed that class oppression was primal over race and gender (Marx and Engels, 1848); Douglas Baynton has explained how disability has often justified oppression towards a variety of social groups (e.g., homosexuality being classified as a mental disorder, women classified as excessively emotional or physically weak) (Baynton, 2001); and Mills has argued for the primacy of race (Mills, 1999). As part of his argument, Charles Mills pushes for the importance of the historical specificity of groups (Mills, 1994). He argues against the Oppression Symmetry Thesis (“symmetry about all oppressions, or at least the Big Three: class, race, gender”) (Mills, 1999), showing that the moral and causal claims are different among the oppressions. Other work acknowledges the racialized and gendered differences to oppression, but argues this does not preclude a unifying theory of class oppression (Sacks, 1989). O’Neill argues there is a tension between idealized notions of justice, which are indifferent to attributes like gender and nationality, and localized notions of justice, which may be insufficiently critical of existing, traditionally endorsed differences such as the household subjugation of women. They argue there are ways to thread this needle by abstracting the social context without idealizing (i.e., to not treat the abstract individual being oppressed as an “ideal” one), and recognizing differences between groups without legitimizing them (i.e., do not accept culturally specific principles) (O’Neill, 1990).
@@ 4. Axis-Level Differences
Next we discuss the axis-level differences between forms of discrimination. Each of these affect both which axes are suitable for which kinds of analysis, as well as which harms should be measured and mitigated for which kinds of axes.
@@ 4.1. Geographic contexts
Machine learning implications: Identity axes should not be taken to be global. Though this may appear obvious, the centrism of American and WEIRD populations is prevalent; prior work found that 84% of analyzed FAccT papers used participants from Western countries, with 63% using participants only on the United States (Septiandri et al., 2023). Other work has pointed out culture-specific distinctions and formulated datasets and fair ML perspectives specifically for the Indian context (Bhatt et al., 2022; Dev et al., 2023; Sambasivan et al., 2021). They explain that the Western-centric focus on race and gender often ends up neglecting subgroups like caste and the context-specificity of religion (e.g., what is a minority religion in one country is a majority one in another) (Sambasivan et al., 2021). For fair ML researchers it can be worth thinking about whether there is a gravitation towards problems with clear racial and gender disparities, to the neglect of other kinds of oppressions.
@@ 4.2. Label availability
Machine learning implications: Not only does data availability dictate which analyses are even permissible, the gradient along the forms of data which are available can also matter. The act of inferring group labels holds with it different normative and empirical implications. In text or tabular data, race and ethnicity can be estimated using BISG, a method which uses surname and geocoded information (Elliott et al., 2009). Specifically for these methods, existing techniques quantify the noise of the racial estimates (McCartan et al., 2024; Lu et al., 2024). In image data, group labels are sometimes visually inferred, with a distinct set of techniques to correct for the noise (Teo et al., 2023). However, for image data there are distinct harms from inferring gender from images (Hamidi et al., 2018; Keyes, 2018). While there are also harms to inferring gender from text (e.g., deadnaming) or tabular data names, they are of a different form than vision due to the harms of visual misgendering and gender performance (Butler, 1990). Inferring race or age from images may also be inaccurate, harmful, and misconstrued (Hanna et al., 2020), but still do not pose the same kinds of harms. Thus, when collecting or inferring labels to use for fair ML, we should consider how each axis has different normative harms, empirical noise, and privacy concerns associated with the label and determine what is appropriate.
@@ 4.3. Legal ramifications
It has been well-studied how the use of attribute-aware methods (i.e., those which take in as input an attribute or proxy attribute label) are in fact sometimes legally impermissible due to antidiscrimination legislation (Ho and Xiang, 2020; Bogen et al., 2020). If the government wishes to use racial classifications, even for cases of reversing historical discrimination, strict scrutiny will apply (ada, 1995). Strict scrutiny is the highest standard of review and requires that distinguishing between races further a “compelling government interest” and be narrowly tailored to achieve a specific interest (Congressional Research Service, 2023). While strict scrutiny applies for suspect classifications such as race, nationality, and religion, a different “intermediate scrutiny” applies to gender (cra, 1976). Compared to strict scrutiny, intermediate scrutiny requires an “important” rather than “compelling” government interest, and the law need only be “substantially related” to the objective rather than “narrowly tailored.” What this means in practice is the government may be more lenient towards the explicit usage of gender in an algorithm compared to that of race (Levinson, 2010). Meanwhile, there is also the notion of “heightened scrutiny,” which is currently legally unclear as to whether it differs from “intermediate scrutiny,” and applies to classifications based on sexual orientation (usw, 2013; smi, 2014). All, however, put burden of proof on government to justify the discrimination.
@@
Additionally, when considering legal discrimination, what counts varies across axes and groups. For example, when age is used as a protected attribute in academic fair machine learning, it is often bucketed into an arbitrary number of categories, or binary based on a threshold. However, legally the Age Discrimination in Employment Act of 1967 (ADEA)222 https://www.eeoc.gov/statutes/age-discrimination-employment-act-1967 is about employment discrimination against people above the age of 40. There are other acts, for example the Age Discrimination Act of 1975 which applies in the domain of federal financial assistance, that applies to other age groups. [^3] https://www.dol.gov/agencies/oasam/regulatory/statutes/age-discrimination-act In fact, the ADEA is actually narrower than disparate impact theory for non-age protected attributes tend to be, permitting otherwise prohibited actions (smi, 2005). In other words, not only is the four-fifths rule not necessarily disparate impact (Watkins et al., 2024), but disparate impact itself might not even be relevant depending on which axis and domain is considered.
@@
As another example of differences in the legal implications of disparities, pay gap discrimination may be covered differently for different groups. While the Equal Pay Act (EPA) 444 https://www.eeoc.gov/equal-paycompensation-discrimination applies only to gender, but not race, Title VII covers both. Wage discrimination claims differ under each because under the EPA, the jobs being compared must be substantially similar, while Title VII does not require that. Additionally, with the EPA the plaintiff does not need to show the employer had discriminatory intent. The burden of proof and damages differ between the two as well, indicating another reason that analyses performed for wage discrimination may differ depending on the attribute being considered.
@@ 5.1. Residual categories
Though the group labels selected for fair ML matter significantly, what is chosen is often simply the convenient choice of binary attributes. For race this is usually Black and White, for gender: men and women, and for disability: disability or not. Each categorization leaves various groups out as the residual categories (i.e., those left out of dominant categorization) (Star and Bowker, 2007). For example, in 2023 a racial categorization of Americans as Black or White would leave out 11% of the population, while a gender categorization of men and women would leave out 1-2% (Brown, 2022); none would be left out for disability because of the way it is defined.
@@
Machine learning implications: The question of how to label individuals to a group has important implications for both training and evaluation. Fairness through awareness broadly describes the category of attribute-aware methods which group labels as input (Dwork et al., 2012). One example is having different thresholds for individuals with different attributes (Hardt et al., 2016). Another is having race-adjusted scores such as in the medical setting (Zink et al., 2024; Burchard et al., 2003). In each case, we need a way to treat the individuals in the residual categories. In terms of evaluation, whereas different kinds of double counting for Multiracial individuals as being part of two groups might be informative, the same kind of double counting of non-binary people into different gender groups could be harmful. On the other hand, certain formulations like “gender minority” may intentionally cluster groups for alliance-building reasons.
@@ 6. Case Study: Chatbot Math Tutoring
At this point, we present a brief case study to make concrete some of the considerations we have discussed so far. Consider a fairness analysis of math tutoring in English through a chatbot for an introduction to algebra course. We may start by identifying which identity axes are relevant, and which harms are salient along each. For gender, we may decide the primary concerns are with respect to ensuring there is no differential treatment, and thus measure for invariance with respect to the student’s gender as well as the content of math word problems. While for race we similarly do not want students of different races to be treated differently, there are two correlated dimensions that could actually warrant differential treatment (Wang et al., 2025), in contrast to enforcing invariance: culture and linguistics. Prior work has shown that culturally and linguistically relevant word problems can make a difference for math education (Driver and Powell, 2017), so we should consider incorporating these findings for tutoring chatbots. Next, when considering age, we might expect many users to be adolescents. This brings up a level of school-appropriateness that would need to be enforced. However, there are also adult learners whom we would want to ensure not to infantilize. Finally, the most relevant axis here is likely learning ability. This is a core topic in education, and covered extensively in the literature (Swanson et al., 2014). In our analysis, we may determine that axes like sexual orientation and marital status are not especially relevant. Taken together, the four axes of gender, race, age, and learning ability are not an exhaustive look at the fairness-related analyses that would need to be considered for this scenario, but an illustrative example of how axis specificity can shape the kinds of harms we measure and mitigate for. We should further consider intersectional interactions if we have specific hypotheses to support them. What we are proposing is in contrast to, for instance, measuring treatment or outcome disparities across all available groups.
@@ 7. Discussion
In this work, we have argued for the specificity of identity axes and groups. This goes against the core values of machine learning, which strive for generality and plug-and-play methods (Birhane et al., 2022). Methods and measurements which generalize across axes help ease adoption, which otherwise serves as a serious roadblock for responsible machine learning. Theories which draw connections between different forms of discrimination can lead to grounded methods, and fruitful collaborations and coalitions across interest groups. There truly are many parts of the machine learning pipeline that are amenable to the substitution of any axis. For disabled or transgender groups whose algorithmic concerns have been historically ignored, something as small as adding a label can now allow the harm to those groups to be measured and potentially mitigated. However, what this framework misses is that specific groups can experience distinct harms. For example, misgendering is a unique harm towards transgender individuals that still be missed (Ovalle et al., 2023, 2024).
@@
One of the greatest fears by computer scientists of incorporating intersectionality into machine learning has been the problem of “exploding groups” where the number of groups to consider exponentially increases. However, incorporating context specificity actually helps to select which axes and groups are needed. In a given domain, one can consult existing regulations to identify legally impermissible practices and relevant protected groups, while also considering historical context to understand which social groups have faced discrimination in the past. In other words, incorporating group specificity can reduce the number of axes and groups studied for a particular harm. At the same time, in cases where greater numbers of groups is a “benefit” to the researcher or practitioner, such as being able to scale a benchmark to be much larger, the exploding groups problem has been leveraged in machine learning to artificially inflate the size of benchmarks. For example, a benchmark might claim over 450,000 unique sentence prompts. However, this scale is only achieved by having around 600 demographic groups and 26 sentence templates multiplied by a number of descriptor terms (Smith et al., 2022). However, as we showed in Fig. 1, this kind of generalization often does not make sense. In this work our goal is to confront fair machine learning’s unyielding pursuit of generality. We do not wish to categorically stop this pursuit, but rather force justification in each instance as to whether generality makes sense. By doing so, we hope to broaden the scope of study to include those harms which are only relevant for one axis or group, treating them just as worthy of concern and science as those harms which are relevant to all axes.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732033isbn: 979-8-4007-1482-5/2025/06ccs: Social and professional topics User characteristicsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Central to most fair machine learning algorithms and measurements are demographic axes, and the groups within the axis. For example, fairness evaluations will measure the difference in outcomes across the demographic axis of gender for the groups of men and women. However, which identity axes or groups usually do not matter to the algorithm or measurement, as the axis is generally left an open variable. In this work, we argue for the importance of specificity in demographic axis. In other words, you cannot necessarily build or measure the fairness of a machine learning system which is fair with respect to race the same way you can for one with respect to gender. While there is increasingly recognition for domain-specific considerations that will affect the relevant definitions of fairness, variability in the axis of identity being targeted remains an often-overlooked dimension.
Fair machine learning has inherited the propensity from machine learning to seek abstractions and generalization, prioritizing methods which are domain-agnostic (Birhane et al., 2022). The ideal model architecture is one which works well across many data distributions. This mentality has led to methods that, for example, treat discrimination as a disparity that occurs between any two social groups. This is not always a bad thing. The ease of implementation often determines whether a quantity is measured at all (Wang et al., 2024), so shoehorning a measure for an overlooked axis (e.g., disability discrimination) into an existing pipeline (e.g., for measuring racial discrimination), can bring more attention than if axis-specific pipelines were to be established in each setting. Similarly, once pipelines for multi-group fairness exist, it seems that so too do those for intersectionality if it can be massaged into the same technical format (Wang et al., 2022; Kong, 2022). Because of the technical convenience of one-size-fits-all fairness, we have generally desired one “fair” algorithm and one “fair” definition across all domains. A common refrain we see in the problem formulation of technical machine learning papers on group fairness is a phrase like “$A$ represents the protected attribute (e.g., race or gender). $A=1$ is the privileged group, and $A\neq 1$ is the unprivileged group” (Mehrabi et al., 2021; Pessach and Shmueli, 2022). But does it matter whether $A$ is race or gender? What about age, or disability?
@@ 2.2. Differences
In considering various forms of oppression, there are debates about the primacy and causal relationships amongst them. Karl Marx believed that class oppression was primal over race and gender (Marx and Engels, 1848); Douglas Baynton has explained how disability has often justified oppression towards a variety of social groups (e.g., homosexuality being classified as a mental disorder, women classified as excessively emotional or physically weak) (Baynton, 2001); and Mills has argued for the primacy of race (Mills, 1999). As part of his argument, Charles Mills pushes for the importance of the historical specificity of groups (Mills, 1994). He argues against the Oppression Symmetry Thesis (“symmetry about all oppressions, or at least the Big Three: class, race, gender”) (Mills, 1999), showing that the moral and causal claims are different among the oppressions. Other work acknowledges the racialized and gendered differences to oppression, but argues this does not preclude a unifying theory of class oppression (Sacks, 1989). O’Neill argues there is a tension between idealized notions of justice, which are indifferent to attributes like gender and nationality, and localized notions of justice, which may be insufficiently critical of existing, traditionally endorsed differences such as the household subjugation of women. They contend there are ways to thread this needle by abstracting the social context without idealizing (i.e., to not treat the abstract individual being oppressed as an “ideal” one), and recognizing differences between groups without legitimizing them (i.e., do not accept culturally specific principles) (O’Neill, 1990; Mills, 2005).
@@ 4. Axis-Level Differences
Next we discuss the axis-level differences between forms of discrimination. Each of these affect both which axes are suitable for which kinds of analysis, as well as which harms should be measured and mitigated for which kinds of axes. For each difference we first provide context then outline the implications for machine learning.
@@ 4.1. Geographic contexts
Machine learning implications: Identity axes should not be taken to be global. Though this may appear obvious, the centrism of American and WEIRD populations is prevalent; prior work found that 84% of analyzed FAccT papers used participants from Western countries, with 63% using participants only in the United States (Septiandri et al., 2023). Other work has pointed out culture-specific distinctions and formulated datasets and fair ML perspectives specifically for the Indian context (Bhatt et al., 2022; Dev et al., 2023; Sambasivan et al., 2021). They explain that the Western-centric focus on race and gender often ends up neglecting subgroups like caste and the context-specificity of religion (e.g., what is a minority religion in one country is a majority one in another) (Sambasivan et al., 2021). For fair ML researchers it can be worth thinking about whether there is a gravitation towards problems with clear racial and gender disparities, to the neglect of other kinds of oppressions.
@@ 4.2. Label availability
Machine learning implications: Not only does data availability dictate which analyses are even permissible, the gradient along the forms of data which are available can also matter. The act of inferring group labels holds with it different normative and empirical implications. In text or tabular data, race and ethnicity can be estimated using BISG, a method which uses surname and geocoded information (Elliott et al., 2009). Specifically for these methods, existing techniques quantify the noise of the racial estimates (McCartan et al., 2024; Lu et al., 2024). In image data, group labels are sometimes visually inferred, with a distinct set of techniques to correct for the noise (Teo et al., 2023). However, for image data there are distinct harms from inferring gender from images (Hamidi et al., 2018; Keyes, 2018). While there are also harms to inferring gender from text (e.g., deadnaming) or tabular data, they are of a different form than vision due to the harms of visual misgendering and gender performance (Butler, 1990). Inferring race or age from images may also be inaccurate, harmful, and misconstrued (Hanna et al., 2020), but still do not pose the same kinds of harms. Thus, when collecting or inferring labels to use for fair ML, we should consider how each axis has different normative harms, empirical noise, and privacy concerns associated with the label and determine what is appropriate.
@@ 4.3. Legal ramifications
It has been well-studied how the use of attribute-aware methods (i.e., those which take in as input an attribute or proxy attribute label) are in fact sometimes legally impermissible due to antidiscrimination legislation (Ho and Xiang, 2020; Bogen et al., 2020). If the government wishes to use racial classifications, even to remedy historical discrimination, strict scrutiny will apply (ada, 1995). Strict scrutiny is the highest standard of judicial review and requires that distinguishing between races further a “compelling government interest” and be narrowly tailored to achieve a specific interest (Congressional Research Service, 2023). While strict scrutiny applies for suspect classifications such as race, nationality, and religion, a different “intermediate scrutiny” applies to gender (cra, 1976). Compared to strict scrutiny, intermediate scrutiny requires an “important” rather than “compelling” government interest, and the law need only be “substantially related” to the objective rather than “narrowly tailored.” What this means in practice is the government may be more lenient towards the explicit usage of gender in an algorithm compared to that of race (Levinson, 2010). A further category, “heightened scrutiny,” applies to sexual orientation,though courts have not clearly distinguished it from intermediate scrutiny yet (usw, 2013; smi, 2014). In all of these cases, however, the burden of proof is on the government to justify the discrimination.
@@
Beyond forms of scrutiny, what counts as legal discrimination also varies across axes and groups. For instance, in fair machine learning research, age is often operationalized by being bucketed into an arbitrary number of categories, or binary based on a threshold. But legally, the Age Discrimination in Employment Act of 1967 (ADEA) [^2] https://www.eeoc.gov/statutes/age-discrimination-employment-act-1967 prohibits employment discrimination only against people above the age of 40. Other laws cover different contexts, e.g., the Age Discrimination Act of 1975 extends protections to additional age groups in the domain of federal financial assistance. [^3] https://www.dol.gov/agencies/oasam/regulatory/statutes/age-discrimination-act In fact, the ADEA is also narrower than other anti-discrimination doctrines: it allows certain practices that would be prohibited under disparate impact theory for race or gender (smi, 2005). In other words, not only is the “four-fifths rule” not necessarily disparate impact (Watkins et al., 2024), but disparate impact itself might not even apply depending on which axis and domain is considered.
@@
Wage discrimination provides another example of how legal implications diverge for different groups. The Equal Pay Act (EPA) 444 https://www.eeoc.gov/equal-paycompensation-discrimination applies only to gender, but not race, while Title VII covers both. Claims also differ in substance: under the EPA, the jobs being compared must be substantially similar, whereas Title VII has no such requirement. Additionally, with the EPA the plaintiff does not need to show the employer had discriminatory intent. The burden of proof and damages differ between the two as well, indicating another reason that analyses performed for wage discrimination may differ depending on the attribute being considered.
@@ 5.1. Residual categories
Though the group labels selected for fair ML matter significantly, what is chosen is often simply the convenient choice of binary attributes. For race this is usually Black and White, for gender: men and women, and for disability: disability or not. Each categorization leaves various groups out of the dominant categorization as the residual categories (Star and Bowker, 2007). For example, in 2023 a racial categorization of Americans as Black or White would leave out 11% of the population, while a gender categorization of men and women would leave out 1-2% (Brown, 2022); none would be left out for disability because of the way it is defined.
@@
Machine learning implications: The question of how to label individuals to a group has important implications for training, prediction, and evaluation. For training, constrained optimization approaches may employ group labels to enforce a fairness constraint. In terms of prediction, fairness through awareness broadly describes the category of attribute-aware methods which group labels as input (Dwork et al., 2012). One example is having different thresholds for individuals with different attributes (Hardt et al., 2016). Another is having race-adjusted scores such as in the medical setting (Zink et al., 2024; Burchard et al., 2003). In each case, we need a way to treat the individuals in the residual categories. And finally for evaluation, whereas different kinds of double counting for Multiracial individuals as being part of two groups might be informative, the same kind of double counting of non-binary people into different gender groups could be harmful. On the other hand, certain formulations like “gender minority” may intentionally cluster groups for alliance-building reasons.
@@ 6. Case Study: Chatbot Math Tutoring
Now that we have completed our presentation of axis-level and group-level differences, we present a brief case study to make concrete some of the considerations we have discussed so far. Consider a fairness analysis of math tutoring in English through a chatbot for an introduction to algebra course. We may start by identifying which identity axes are relevant, and which harms are salient along each. For gender, we may decide the primary concerns are with respect to ensuring there is no differential treatment, and thus measure for invariance with respect to the student’s gender as well as the content of math word problems. While for race we similarly do not want students of different races to be treated differently, there are two correlated dimensions that could actually warrant differential treatment (Wang et al., 2025), in contrast to enforcing invariance: culture and linguistics. Prior work has shown that culturally and linguistically relevant word problems can make a difference for math education (Driver and Powell, 2017), so we should consider incorporating these findings for tutoring chatbots. Next, when considering age, we might expect many users to be adolescents. This brings up a level of school-appropriateness that would need to be enforced. However, there are also adult learners whom we would want to ensure not to infantilize. Finally, the most relevant axis here is likely learning ability. This is a core topic in education, and covered extensively in the literature (Swanson et al., 2014). In our analysis, we may determine that axes like sexual orientation and marital status are not especially relevant. Taken together, the four axes of gender, race, age, and learning ability are not an exhaustive look at the fairness-related analyses that would need to be considered for this scenario, but an illustrative example of how axis specificity can shape the kinds of harms we measure and mitigate for. We should further consider intersectional interactions if we have specific hypotheses to support them. What we are proposing is in contrast to, for instance, measuring treatment or outcome disparities across all available groups.
@@ 7. Discussion
In this work, we have argued for the specificity of identity axes and groups. This goes against the core values of machine learning, which strive for generality and plug-and-play methods (Birhane et al., 2022). Methods and measurements which generalize across axes help ease adoption, which otherwise serves as a serious roadblock for responsible machine learning. Theories which draw connections between different forms of discrimination can lead to grounded methods, and fruitful collaborations and coalitions across interest groups. There truly are many parts of the machine learning pipeline that are amenable to the substitution of any axis. For disabled or transgender groups whose algorithmic concerns have been historically ignored, something as small as adding a label can now allow the harm to those groups to be measured and potentially mitigated. However, what this framework misses is that specific groups can experience distinct harms. For example, misgendering is a unique harm towards transgender individuals that will be missed (Ovalle et al., 2023, 2024).
@@
One of the greatest fears by computer scientists of incorporating intersectionality into machine learning has been the problem of “exploding groups” where the number of groups to consider exponentially increases. However, incorporating context specificity actually helps to select which axes and groups are needed. In a given domain, one can consider historical context to understand which social groups have faced discrimination in the past, as well as consult existing regulations to identify legally impermissible practices and relevant protected groups. In other words, incorporating group specificity can reduce the number of axes and groups studied for a particular harm. At the same time, in cases where greater numbers of groups is a “benefit” to the researcher or practitioner, such as being able to scale a benchmark to be much larger, the exploding groups problem has been leveraged in machine learning to artificially inflate the size of benchmarks. For example, a benchmark might claim over 450,000 unique sentence prompts. However, this scale is only achieved by having around 600 demographic groups and 26 sentence templates multiplied by a number of descriptor terms (Smith et al., 2022). However, as we showed in Fig. 1, this kind of generalization often does not make sense. In this work our goal is to confront fair machine learning’s unyielding pursuit of generality. We do not wish to categorically stop this pursuit, but rather force justification in each instance as to whether generality makes sense. By doing so, we hope to broaden the scope of study to include those harms which are only relevant for one axis or group, treating them just as worthy of concern and science as those harms which are relevant to all axes.